OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of each should he use? Let p = pounds of peanuts and let c = pounds of cashews. Write a system of equations that could be used to solve the problem.

OpenStudy (anonymous):

C+P=10 You can pick any combination of pounds for cashews and peanuts but it has to be a 10 pound mixture so that's why we set it up to 10

OpenStudy (anonymous):

ook thx now that makes since

OpenStudy (anonymous):

p+c = 10 .....(i) {4p+6.50c}/10 = 4.75 .......(ii) solve these equations and u will get the answer. p=7 and c =3 therefore he should use 7 pounds of peanuts and 3 pounds of cashews